OpenStudy (anonymous):

Which fraction expresses -10/45 in lowest terms? A.-1/3 B.-2/9 C.-1/9

OpenStudy (anonymous):

What do you think it is?

OpenStudy (whpalmer4):

Can you write a multiplication of prime numbers (2,3,5,7,11,13,17,19,etc.) that gives you 10? How about one that gives you 45? If any numbers appear in both multiplications, you can divide both the top and the bottom of the fraction by that number to get a simpler fraction.

OpenStudy (anonymous):

-2/9?

OpenStudy (whpalmer4):

\[-\frac{2}{9} = -\frac{2*5}{9*5} = -\frac{10}{45}\checkmark\]

OpenStudy (anonymous):

@diamondrowland22 You are correct.

OpenStudy (anonymous):

okay thanks